Greenhab Report – April 20th

Crew 211 GreenHab Report – 20-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with the door open

Shade cloth: 80%

Average temperatures:

Low: 14°C

High: 43°C

Hours of supplemental light: 0

Daily water usage for crops: 2 gallons

Water in Blue Tank: 125 gallons

Time(s) of watering for crops: 19:30

Changes to crops: Crew 211 has just arrived today

Narrative: Crew 211 has just arrived today

Harvest: None

Support/supplies needed: None

Operations Report – April 20th

Crew 211 Operations Report 20-APR-2019

SOL: /

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 12
On (last night): 7 pm
Off (this morning): 7 am

Solar— SOC 79 % at 7 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 475 gallons

Auxiliary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 125 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 9 units

Toilet tank emptied: no

Deimos rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes
Hours: 51.1h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: no
Hours: 87.6h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 56.8h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Curiosity rover used: no
Hours: 89.1h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? Yes, the trip from Grand Junction to MDRS

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: No operation, Chloe will send the first GreenHab report today

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: None, thank you.

Sol Summary – April 20th

Crew 211 Sol Summary Report 20-APR-2019

Sol: /

Summary Title: Discovering the MDRS

Author’s name: Eléonore Lieffrig

Mission Status: Nominal

Sol Activity Summary:

Today was a day of traveling and discovering the MDRS. Five members of the crew arrived at the station at about 1:30 pm. As Chloé, Julien and I discovered our new home, Carl-Henrik and Simon went back to Grand Junction in order to take the rest of the crew and give back our rental car. At the moment (7 pm), they are still on their way.

We took our marks here, cleaned a bit the Hab, chose our bedrooms and got to know the place. Shannon gave us a brief tour of the station, but a longer training is planned with the whole crew tomorrow morning at 10 am.

Tonight, we will read the Hab manual, welcome the rest of the team, and cook ourselves our first tasty Martian dinner.

For the crew 211,

Eléonore Lieffrig, astronomer

Anomalies in work: No problems.

Weather: Warm and windy

Crew Physical Status: Good.

EVA: No EVA today.

Reports to be filed: GreenHab Report, Operations report, Sol Summary Report

Support Requested: None

Sol Summary – April 19th

Crew 210 Sol Summary Report 19-APR-2019

Sol: 6

Summary Title: Writing and preparing for crew handover

Author’s name: Paul Sokoloff

Mission Status: Nominal

Sol Activity Summary: Today the crew prepared to depart MDRS and leave it to our successors on Crew 211. We spent the morning cleaning the station, inventorying the food, and getting the spaces ship-shape for handover.

In the afternoon we continued to process plant samples and salinity measurements in the Science Dome, and collaboratively wrote a white paper outlining our multi-year biodiversity program, while Mike Irvine prepared an in-depth outreach proposal.

We will continue to prepare for our imminent departure this evening.

Look Ahead Plan: We plan to leave MDRS at 0800 tomorrow to head to Grand Junction, where we plan on handing over the crew car to Crew 211 by 1200.

Thank you to Mission Support, the Mars Society, and the management of MDRS for a great week in Utah; the five of us really appreciate your support.

Anomalies in work: No problems.

Weather: Hot and sunny.

Crew Physical Status: Good.

EVA: No EVAs today.

Reports to be filed: GreenHab Report, Operations Report, Daily Photo Report, Mission Summary, End of Rotation Food Inventory, End of Rotation Checklist (by tomorrow am).

Support Requested: None

Greenhab Report – April 19th

GreenHab Officer: David Murray

Environmental control: Ambient with fan 0930-1900

Shade Cloth: 80%

Low Temperature: 15 C

High Temperature: 37 C

Relative humidity: 15%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 34 L total, 8 L fertilizer

Water in Blue Tank: refilled with about 85 gallons of water

Time(s) of Watering: 1100 and 1850

Changes to Crops: Tomatoes are stressed by the heat (curling, drooping, yellowing of basal foliage)

Narrative: It was very hot out today and the tomatoes didn’t like that. The other plants look healthy. I fertilized again because not everything was fertilized the first time and the first application was light. The fan was on from 0930 to 1900 but was turned off while filming a GreenHab video.

Tonight was my last night as the GreenHab officer because Crew 210 departs for Grand Junction at 0800 tomorrow morning.

Harvest: None

Operations Report – April 19th

Crew 210 Operations Report 19-APR-2019

SOL: 7

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 12

On (last night): 7 pm

Off (this morning): 7 am

Solar— SOC 75 % at 7 pm (so turned generator on)

Diesel Reading – 34 %

Propane Reading – 49%

Ethanol Free Gasoline – 4 gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 425 gallons

Auxiliary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 150 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: not read

Toilet tank emptied: no

Deimos rover used: Still non-operational

Hours:
Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, no

Hours: Did not record

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours: did not check

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: not recorded

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours: did not record

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) all

Reason for use: moving them back over to Shannon’s house

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? Yes, a trip to town

General notes and comments: Submersible pump in the transfer water tank is working

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Cleaning

Summary of GreenHab operations: David will send a full report today.

Summary of ScienceDome operations: Finishing up lab work and cleaning

Summary of RAM operations: Cleaning

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing tonight,
thank you.

Science Report – April 18th

Science Report 18 April 2019
Crew 210
Submitted by Paul Sokoloff

Today we continued to process the soil samples taken at each of the 12 transects we conducted over the last few days. In the Science Dome, we measured the pH and salinity and characterized the texture of the soil samples. These data will be used to characterize the ecotones of the deserts around MDRS in an upcoming paper.

We continued to collect new plant specimens opportunistically, including an annual invasive grass, (Annual Wheatgrass – Eremopyrum triticeum) newly recorded for MDRS.

We plan on continuing our lab work and wrapping up data recording for the plant project tomorrow.

Greenhab Report – April 18th

GreenHab Officer: David Murray

Environmental control: Ambient, fan on from 0930 to 1900; the door closed.

Shade Cloth: 80%

Low Temperature: 24 C

High Temperature: 33 C

Relative humidity: 15%

Hours of Supplemental Light: 0

Daily Water Usage for Crops: 38 L

Water in Blue Tank: Refilling tank in about one hour

Time(s) of Watering: 1030

Changes to Crops: The tomatoes didn’t like the heat today but they’re alright. Tomato starts haven’t shown much growth but everything else is responding very well to the fertilizer.

Narrative: The door was left closed with the fan on high from 0930 to 1900 with passive exhaust as part of an experiment. The internal temperature stayed warm but fluctuated less drastically than it usually does with the door open.

Although the high was above 30 degrees C, I think the average temperature was right around perfect for all of the crops but the tomatoes (and possibly the beans).

Harvested basil, sage, carrot, and parsley.

Harvest: 7 g basil, 7 g parsley, 15 g carrot, 1 g sage